solve for x
4x(x-4)-5x(x-3)=0
thank you

Answer by John10(297)

solve for x

4x(x-4)-5x(x-3)=0
Multiply on the left side using FOIL method
4x^2 - 16x - 5x^2 + 15x = 0
Collect the like terms:
(4x^2 -5x^2) + (-16x +15x) = 0
-x^2 -x = 0
Factor it out
-x(x+1)=0
Use the zero product property to solve for x:
-x = 0 OR x + 1 = 0
You can get it from here. Remember to check your answers!
Thanks! John10